OpenStudy (anonymous):

limit of 3(1-cosx)/x as x approaches 0

OpenStudy (anonymous):

do u know the l hopital rule?

OpenStudy (anonymous):

have u tried it yet?

OpenStudy (anonymous):

i am gong to make a guess that you have not got to derivatives yet, let alone l'hopital's rule am i right?

OpenStudy (anonymous):

Yeah you're right

OpenStudy (anonymous):

then here is what you need to know: \[\lim_{x\to 0}\frac{\cos(x)-1}{x}=0\]

OpenStudy (anonymous):

the 3 out front is not important because \(3\times 0=0\)

OpenStudy (anonymous):

Well alright. Thanks!

OpenStudy (anonymous):

the other one you need to know is \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\] you will need these to find the derivatives of sine and cosine they are usually proved in a not very convincing way by geometry and an appeal to something called the "squeeze theorem" but you don't need to prove it, just try to remember it

OpenStudy (anonymous):

yw